OpenStudy (anonymous):

oops try this determine whether the given quadratic has a min or max value then find the min or max value and determine where it occurs f(x)=4x^2+2x+2

OpenStudy (saifoo.khan):

Compare it with: ax^2 + bx + c = 0 the values of a is +4. this means that we will have a min point.

OpenStudy (anonymous):

min value is 7/4

OpenStudy (saifoo.khan):

To determine where it occurs, use -b/2a to find it's "x" cordinate. the value of b is 2 and a is 4. \[-\frac{b}{2a} \to -\frac{2}{2(4)} \to -\frac{2}{8} \to -\frac14\]

OpenStudy (saifoo.khan):

Now insert this value in the equation to get the "y" cordinate.